# Concept Bottleneck Language Models For protein design

### v1

@@ 2.1 Architecture
The overall architecture is shown in Figure 2. We start with a standard BERT-based masked language model (MLM) (Devlin et al., 2018), which is commonly used in many protein language models (Rives et al., 2019; Lin et al., 2022; Hayes et al., 2024; Frey et al., 2024). A mask token [MASK] is randomly applied to a subset of tokens in the sequence $x$ resulting in as masked version of the sequence . After masking, ${x}$ is passed into a multi-head self-attention transformer encoder, and the output of the encoder is denoted as $\mathbf{H}=\text{TransformerEncoder}({x})$, where $\mathbf{H}=[\mathbf{h}_{0},\dotsc,\mathbf{h}_{s}]$. We introduce three main architectural changes in CB-LM: (a) concept bottleneck module (CB-module), (b) orthogonality network, and (c) linear decoder.
@@ 3.1 Model quality
Models 24M 150M
LBSTER (No conditioning) 12.17 9.92
ESM2 (No conditioning) - 9.58
C-pLM 11.65 10.57
CC-pLM 12.37 11.24
CB-pLM 11.57 10.49
Table 1: Validation perplexity.
Model quality is evaluated using perplexity, which measures how well the model predicts a token based on its sequence. Perplexity ranges from 1 (perfect model) to the vocabulary size (33 for the pLMs vocabulary (Lin et al., 2022)), with higher values indicating random predictions. We assessed performance on a held-out validation set. Table 1 presents the validation perplexity of different models, where LBSTER (Frey et al., 2024) and ESM2(Lin et al., 2022) are standard masked language models without conditioning. C-pLM and CC-pLM use ground-truth concepts as inputs; since we lack actual values for all concepts and can only compute Biopython concepts from the sequence, these models were given Biopython concepts with the rest set as missing values. All models learned the training distribution well, with similar-sized models showing comparable perplexity. Notably, CB-pLM achieves higher perplexity among conditioning models, likely because it predicts concepts rather than requiring additional input information. CB-pLM matches state-of-the-art open-source protein masked language models performance, even surpassing them in the 24M parameter models. These results are promising, as CB-pLM, despite its constraints, shows no significant performance drop compared to unconstrained models. Scaling: To verify that our proposed architecture can scale to larger models we train a 650M and 3B parameter CB-pLM with perplexities 9.07 and 9.64 respectively.
@@ 4.2.1 Experiments
Baselines We consider a wide variety of representative approaches to guided and unguided sequence generation: (a) Unguided Discrete Walk-Jump Sampling (WJS) (Frey et al., 2023) energy- and score-based model with single-step denoising. (b) Implicitly guided PropEn (Tagasovska et al., 2024) an encoder-decoder architecture that is implicitly trained to optimize a property of interest. (c) Explicitly guided LaMBO-2 (Gruver et al., 2024) a classifier-guided discrete diffusion model. Additional details about baselines and training procedures are available in Appendix C.2.
Results We intervene on the GRAVY concept for generated samples for each candidate protein design model, with the constraint that generated samples should be at most an edit distance of 5 away from the Siltuximab sequence. In Figure 7(a), we show the distribution of designs from each model type. The CB-pLM comes second to LaMBO-2 in terms of the degree of concept distribution shift induced in the generated output, while preserving the naturalness of the protein (Figure 7(b)). Our findings suggests that a general-purpose CB-pLM trained to learn over 700 concepts delivers comparable results to state-of-the-art conditioning approaches while preserving the naturalness of the protein.
@@ 5.2 Debugging
For large models trained on natural language, it is difficult to identify the inputs on which a model will fail;—exhibit undesirable behavior—a challenge now known as LLM red teaming Perez et al. (2022). Such inscrutability is compounded in protein language models because model failure cannot simply detected from inspecting the model’s output. The linear layer, which maps concept representations to an output embedding, helps provide domain expert with ability to identify model failures. For example, by inspecting the weights of the linear layer, we can detect how the output of a model can be controlled by a given concept.
@@ 6 Discussion and Conclusion
Discussion
In this paper, we applied our proposed architecture to masked language models for proteins. This architecture can be adapted to any domain, such as NLP, by simply changing the training dataset, vocabulary, and concepts while keeping the model unchanged. It can also be adapted to other language models, such as autoregressive models, as shown in Figure 24 in Appendix E.2. Our architecture supports the addition of new concepts; given a new dataset, we can fine-tune CB-LM with a new concept by copying weights from the original model, with only the new concept embedding, classifier head, and linear decoder being randomly initialized. The new model is then trained to match the old model’s output for existing concepts while aligning with the ground truth values for the new concept. In addition, CB-LM can be used to generate sequences with unseen concepts or novel combinations, as long as they can be expressed as functions of existing ones. We demonstrate this capability in a toy task detailed in Appendix C.3.2.
@@ A.1 Loss Function
    $\mathcal{L}_{\mathrm{orth}}=\sum_{j\in B}{\frac{\sum_{i=1}^{i=k}\big{|}\langle w_{i}\,,\,w_{k+1}\rangle\big{|}}{\sum_{i=1}^{i=k}1}}$ (1)
@@ C.1.1 Single concept intervention
We investigate how the concept distribution shits along with the naturalness of the protein after intervention. Ideally, we would want the concept distribution to shift in the direction of the intervention while maintaining the naturalness of the protein; we plot $\Delta$ Concept/Naturalness distribution (the difference between the original sample and the generated sample for the concept value on $x$-axis and naturalness on $y$-axis).
@@
In this experiment, we mask the most influential tokens for each concept as previously described and intervene on the value of that concept. We then calculate the value of all 14 concepts for the generated sequence, and see how intervening on one concept effects all the other concepts. The results are reported in Figures 17(b), 17(c) and 17(d) for different models. In this figure we report a version of accuracy, where the $j$-th element of the $i$-th row represents the (fraction of times concept $j$ changed in the same direction as the intervention when intervening on the $i$-th concept) - (fraction of times concept $j$ changed in the opposite direction of the intervention when intervening on the $i$-th concept). This can be seen as a binary version of correlation. We also compare this with the actual correlation between these concept values on existing proteins in our dataset shown in Figure 17(a). We can see the correlation matrix of interventions on our CBM has many similarities to the correlations between the concept in the real world, especially on pairs with high correlation, reflecting that our concept bottleneck model has picked up on these natural correlations. In contrast, the other models have a quite different correlation pattern.
@@ D.2 Understanding concept/output relationships
To understand the model’s ability to infer biophysical properties, we visualized the weights from the final layer for each concept in relation to each amino acid. Our findings indicate that the model successfully learns several key biophysical relationships as defined in the BioPython library (Figures 21, 22 and 23).
@@
  • For both charge at pH 6 and pH 7, acidic amino acids (D, E) have negative weights, while basic amino acids (R, K) have positive weights. The difference in charge between pH 6 and pH 7 reflects the biophysical properties of H , as it contains an imidazole side chain with a pKa of 6.0 and contributes more towards positive charge at lower pH (Figures 22(a), 22(b)).
@@
  • Helix fraction, sheet structure fraction, and turn structure fraction are related concepts, are defined by fractions of I, L, F, W, Y for helix, A, L, M, E for sheet and N, S, G, P. While the weights do not closely match that of the BioPython definitions, we note that the weights are nonetheless consistent with amino acids that often have special propensity to be part of helical structures, such as A, M, L, E, K (Leiro et al., 2017) (Figures 23(a), 23(b) and 23(c)).
@@
  • Molar extinction coefficient, which quantifies absorption of a protein at 280 nm, emphasizes the contributions of W. Likewise, the reduced vs oxidized terms of the molar extinction parameter differ only in how the oxidation state of C is considered, which is also visible in our model weights (Figures 23(f), and 23(e)).

### v2

@@ 2.1 Architecture
The overall architecture is shown in Figure 2. We start with a standard BERT-based masked language model (MLM) (Devlin et al., 2018), which is commonly used in many protein language models (Rives et al., 2019; Lin et al., 2022; Hayes et al., 2024; Frey et al., 2024). A mask token [MASK] is randomly applied to a subset of tokens in the sequence $x$ resulting in a masked version of the sequence. After masking, ${x}$ is passed into a multi-head self-attention transformer encoder, and the output of the encoder is denoted as $\mathbf{H}=\text{TransformerEncoder}({x})$, where $\mathbf{H}=[\mathbf{h}_{0},\dotsc,\mathbf{h}_{s}]$. We introduce three main architectural changes in CB-LM: (a) concept bottleneck module (CB-module), (b) orthogonality network, and (c) linear decoder.
@@ 3.1 Model quality
Model quality is evaluated using perplexity, which measures how well the model predicts a token based on its sequence. Perplexity ranges from 1 (indicating a perfect model) to the vocabulary size (33 for the pLMs vocabulary (Lin et al., 2022)), with higher values indicating more random predictions. To assess performance, we used 10,000 randomly sampled antibodies from the publicly available Mason dataset Mason et al. (2021). Table 1 shows the perplexity of different models on this dataset. LBSTER (Frey et al., 2024) and ESM2 (Lin et al., 2022) are standard masked language models that do not require additional tags during inference and do not support conditioning. Conditioning baselines C-pLM and CC-pLM require ground-truth concepts as inputs during inference; they were given Biopython concepts for sequences, while other concepts were set as missing values. All models learned the training distribution well, with similar-sized models showing comparable perplexity. CB-pLM has a perplexity comparable to conditional models, even though C-pLM and CC-pLM received additional tags as input, while CB-pLM did not. Notably, CB-pLM often achieves higher perplexity than state-of-the-art open-source protein-masked language models of similar size. These results are promising, as CB-pLM, despite its constraints, shows no significant performance drop compared to unconstrained models; in contrast, forcing the model to learn bio-physical concepts might result in a better model overall.
Models Tags Cond. 24M 150M 650M 3B
LBSTER No No 8.20 2.97 - -
ESM2 No No - 4.84 3.41 3.01
C-pLM Yes Yes 4.57 2.62 - -
CC-pLM Yes Yes 4.07 2.74 - -
CB-pLM No Yes 4.29 3.20 2.48 2.50
Table 1: Mason dataset perplexity.
@@ 4.2.1 Experiments
Baselines We consider a wide variety of representative approaches to guided and unguided sequence generation: (a) Unguided Discrete Walk-Jump Sampling (WJS) (Frey et al., 2023) energy- and score-based model with single-step denoising and ESM2 (Lin et al., 2022) a large one-source protein language model. (c) Implicitly guided PropEn (Tagasovska et al., 2024) an encoder-decoder architecture that is implicitly trained to optimize a property of interest. (d) Explicitly guided LaMBO-2 (Gruver et al., 2024) a classifier-guided discrete diffusion model. (e) Hydrophilic Resample a non-deep learning baseline, where residues are randomly resampled from a set of known hydrophilic residues (N, C, Q, G, S, T, Y) (Aftabuddin & Kundu, 2007). Additional details about baselines and training procedures are available in Appendix C.2.
Results We intervene on the GRAVY concept for generated samples for each candidate protein design model, with the constraint that generated samples should be at most an edit distance of 5 away from the Siltuximab sequence. In Figure 7(a), we show the distribution of designs from each model type. The CB-pLM comes second to LaMBO-2 in terms of the degree of concept distribution shift induced in the generated output, while preserving the naturalness of the protein (Figure 7(b)). trained to learn over 700 concepts delivers comparable results to state-of-the-art conditioning approaches while preserving the naturalness of the protein.
@@ 5.1 Interpretability
CB-pLM offers two types of interpretability: Local and Global. Local Interpretability allows the model to indicate which concepts it relies on for predictions and identify the input tokens that influence its output, achieved through our concept-bottleneck module and feature masking and regulation scheme. Global Interpretability provides insights into the relationships between predefined concepts and various tokens, facilitated by the linear decoder. This section explores CB-pLM’s global interpretability and verifies its alignment with human knowledge.
@@ 5.2 Debugging
Model debugging involves detecting and fixing errors during training or testing. Identifying the root cause of an error–such as issues with optimization, initialization, training data, or function class–can be challenging and often relies on manual effort or retraining. CB-LM introduces mechanisms to attribute errors to human-understandable concepts, enabling us to answer key questions: 1) Did the model learn an important feature of interest? 2) Is the model relying on a spurious feature? 3) In which situations is the model likely to fail? 4) How can we correct unwanted correlations? These questions can be addressed by inspecting the validation loss of different concepts, the weights of the linear layer, and using the CB layer for interventions.
@@ 6 Discussion and Conclusion
Discussion and Future Work.
In this paper, we applied our proposed architecture to masked language models for proteins. This architecture can be adapted to any domain, such as NLP, by changing the training dataset, vocabulary, and concepts while keeping the model unchanged. It can also be extended to other language models, such as autoregressive models (Figure 29 in Appendix E.2). In principle, CB-LM can be used to generate sequences with unseen concepts or novel combinations, as long as they can be expressed as functions of existing ones. We demonstrate this capability in a toy task detailed in Appendix C.3.2. Testing this capability empirically on proteins is beyond the scope of the current paper and is left for future work. Our method is unlikely to enable the development of weaponizable molecules due to limited biophysical understanding of biological functions and physiochemical properties. However, we stress the importance of biosecurity and recommend consulting regulatory agencies and ethical guidelines to mitigate potential risks. C.3.2.
@@ Conclusion
Acknowledgement
The authors acknowledge the entire Prescient Design team, the Antibody Engineering department at Genentech, and NVIDIA for providing helpful discussions and input that contributed to the research results reported within this paper. We would especially like to acknowledge Greg Zynda, Bruno Trentini, and Anthony Costa from NVIDIA.
@@ A.1 Loss Function
    $\mathcal{L}_{\mathrm{orth}}=\sum_{j\in B}{\frac{\sum_{i=0}^{i=s}\big{|}\langle z\,,\,\widetilde{h}_{i}\rangle\big{|}}{\sum_{i=0}^{i=s}1}}$ (1)
@@ C.1.1 Single concept intervention
We investigate how the concept distribution shifts along with the naturalness of the protein after intervention. Ideally, we would want the concept distribution to shift in the direction of the intervention while maintaining the naturalness of the protein; we plot $\Delta$ Concept/Naturalness distribution (the difference between the original sample and the generated sample for the concept value on $x$-axis and naturalness on $y$-axis).
@@
Masking percentage
In this experiment, we investigate the impact of varying the percentage of masking on the performance ranking of different models. We repeat the experiment described in Algorithm 1, but with 25% masking instead of 5%, using 24M parameter models. Our findings show that the performance ranking of the models remains consistent, with CB-pLM significantly outperforming both C-pLM and CC-pLM (Figures 17,18 and 19).
(a) Intervention accuracy.
(b) Average change in concept value after intervention.
Figure 17: Average concept intervention accuracy and effectiveness when masking 25%.
Figure 18: Shift in protein naturalness/concept value for positive interventions when masking 25%.
Figure 19: Shift in protein naturalness/concept value for negative interventions when masking 25%.
@@
In this experiment, we mask the most influential tokens for each concept as previously described and intervene on the value of that concept. We then calculate the value of all 14 concepts for the generated sequence, and see how intervening on one concept effects all the other concepts. The results are reported in Figures 20(b), 20(c) and 20(d) for different models. In this figure we report a version of accuracy, where the $j$-th element of the $i$-th row represents the (fraction of times concept $j$ changed in the same direction as the intervention when intervening on the $i$-th concept) - (fraction of times concept $j$ changed in the opposite direction of the intervention when intervening on the $i$-th concept). This can be seen as a binary version of correlation. We also compare this with the actual correlation between these concept values on existing proteins in our dataset shown in Figure 20(a). We can see the correlation matrix of interventions on our CBM has many similarities to the correlations between the concept in the real world, especially on pairs with high correlation, reflecting that our concept bottleneck model has picked up on these natural correlations. In contrast, the other models have a quite different correlation pattern.
@@ C.2.1 Baselines
  • ESM2 (Lin et al., 2022) an open source 150M parameter protein language model, pretrained on UniRef50. Designs are sampled from the model’s logits.
  • Hydrophilic Resample a non-deep learning baseline, where residues are randomly resampled from a set of known hydrophilic residues (N, C, Q, G, S, T, Y) (Aftabuddin & Kundu, 2007).
C.2.2 Assessing the naturalness of the designs
To evaluate the naturalness of designs generated by various models, we folded the designs using ABodyBuilder2 (Abanades et al., 2023) and analyzed different protein surface properties with the Therapeutic Antibody Profiler (TAP) Raybould & Deane (2022), which utilizes physics-based computations. Our objective is to achieve minimal deviations from the original antibody structure. Figure 22 presents the TAP scores of different designs, with the dotted line indicating the original Siltuximab value. The results reveal that the designs produced by CB-pLM and WJS exhibit the least disruption, whereas those from ESM2 and Hydrophilic Resample show the most disruption. The TAP metrics validate that our CB-pLM designs effectively preserve naturalness.
Figure 22: TAP metrics for Siltuximab redesigns.
Figure 23: Percentages of redesigns within the naturalness threshold.
Filtering Designs
Out of the designs produced by different models, we have filtered designs with TAP values that are with 5% proximity to Siltuximab; the percentage of designs that have passed this filter from each model is shown in Figure 23.
@@ D.2 Understanding concept/output relationships
To understand the model’s ability to infer biophysical properties, we visualized the weights from the final layer for each concept in relation to each amino acid. Our findings indicate that the model successfully learns several key biophysical relationships as defined in the BioPython library (Figures 26, 27 and 28).
@@
  • For both charge at pH 6 and pH 7, acidic amino acids (D, E) have negative weights, while basic amino acids (R, K) have positive weights. The difference in charge between pH 6 and pH 7 reflects the biophysical properties of H , as it contains an imidazole side chain with a pKa of 6.0 and contributes more towards positive charge at lower pH (Figures 27(a), 27(b)).
@@
  • Helix fraction, sheet structure fraction, and turn structure fraction are related concepts, are defined by fractions of I, L, F, W, Y for helix, A, L, M, E for sheet and N, S, G, P. While the weights do not closely match that of the BioPython definitions, we note that the weights are nonetheless consistent with amino acids that often have special propensity to be part of helical structures, such as A, M, L, E, K (Leiro et al., 2017) (Figures 28(a), 28(b) and 28(c)).
@@
  • Molar extinction coefficient, which quantifies absorption of a protein at 280 nm, emphasizes the contributions of W. Likewise, the reduced vs oxidized terms of the molar extinction parameter differ only in how the oxidation state of C is considered, which is also visible in our model weights (Figures 28(f), and 28(e)).